A community health system’s nurse team consists of 12 nurses working in the local community. It takes a nurse 1.5 hour to complete one patient visit (including travel times and breaks).

a.  What is the capacity of the nurse team over the course of a 11-hour work day? Round your answer to the nearest whole number. For example, if your answer is 8.654, input 9.

Your answer is .

b.  Assuming the demand for the nurses is 80 patients per day, what is the utilization of the nurse team? Round "Utilization" to a whole percent and ignore the percentage sign. For example, if your answer is 80.32% or 0.8032, input 80.  

Your answer is .

c.  Assuming the demand for the nurses is 80 patients per day, what is the cycle time in minutes? "Cycle Time" to 2 decimals. For example, if your answer is 4.5631, input 4.56.

Your answer is .
Answer #1

Answer a: capacity of the nurse team over the course of an 11-hour work day = capacity of a nurse*no. of nurse*11 hours = (1/1.5)*12*11 = 88 patients

Answer b: demand is less than capacity, therefore demand is actual output

Utilization= Actual out put/capacity = 80/88 = 0.909 or 91%

Answer c: output= 80

11 hour work days

therefore, cycle time= 11*60/80= 8.25 mins per patient

Please ask, if you have any doubts through the comment section. Do rate the answer